### v1

# Verification and Validation for Trustworthy Scientific Machine Learning

Abstract
Scientific machine learning (SciML) models are transforming many scientific disciplines. However, the development of good modeling practices to increase the trustworthiness of SciML has lagged behind its application, limiting its potential impact. The goal of this paper is to start a discussion on establishing consensus-based good practices for predictive SciML. We identify key challenges in applying existing computational science and engineering guidelines, such as verification and validation protocols, and provide recommendations to address these challenges. Our discussion focuses on predictive SciML, which uses machine learning models to learn, improve, and accelerate numerical simulations of physical systems. While centered on predictive applications, our 16 recommendations aim to help researchers conduct and document their modeling processes rigorously across all SciML domains.
1 Introduction
Scientific machine learning (SciML) integrates machine learning (ML) into scientific workflows to enhance system simulation and analysis, with an emphasis on computational modeling of physical systems. This field emerged from Department of Energy workshops and initiatives starting in 2018, which also identified the need to increase “the scale, rigor, robustness, and reliability of SciML necessary for routine use in science and engineering applications” [5]. The field’s subsequent growth through funding initiatives, conference themes, and high-profile publications stems from its ability to unite ML’s predictive power with the domain knowledge and mathematical rigor of computational science and engineering (CSE). However, this surge in SciML development has outpaced good practices and reporting standards for building trust [66, 51, 109, 117].
SciML models must demonstrate trustworthiness to be safe and useful [44]. Organizational and computational trust definitions [92, 106] inform our criteria for trustworthy SciML: competence in basic performance, reliability across conditions, transparency about processes and limitations, and alignment with scientific objectives. These criteria span technical attributes (correctness, reliability, safety) and human-centric qualities (comprehensibility, transparency).111The notion of trustworthiness in SciML differs from “trustworthy AI” in socio-technical systems. While trustworthy AI addresses societal impacts, ethical considerations, and human behavioral aspects, trustworthiness in SciML focuses on mathematical rigor, physical consistency, and computational reliability in scientific and engineering applications. Trustworthy SciML performs consistently across operating conditions while providing insight into decision-making and adhering to scientific principles. The model’s decision process must align with prior information (intrinsic trustworthiness) and generalize to unseen cases (extrinsic trustworthiness) [44].
Rigorous modeling practices require sustained effort from the scientific community. CSE has established standards over decades for model reporting and assessment, including verification and validation, across biology [80], earth sciences [79, 45], and engineering [95, 84, 89, 1]. These standards help users and decision-makers document limitations, uncertainties, and modeling choices to prevent resource misallocation and maintain model credibility. Similar SciML efforts have only recently begun [51, 109, 117]. Developers and users of SciML must establish best practices for the development, reporting, and evaluation of new SciML paradigms.
This paper initiates a dialogue toward consensus-based practices for predictive SciML—the use of ML models to learn, improve, or accelerate physical system predictions. Rather than creating a rigid checklist, we guide transparent modeling processes by adapting existing CSE verification and validation (V&V) standards [1, 4, 72] while addressing SciML’s unique challenges. Our guidance applies to both the development and deployment in support of a scientific claim and new algorithm development.
The remainder of the paper is structured as follows. Section 2 defines our scope, compares CSE and SciML development processes, and presents our four-component framework for SciML model development. Sections 3–6 detail these components and develop the recommendations. The resulting 16 recommendations are listed below for quick reference. Section 7 summarizes the key messages.
2 Background
While SciML is becoming increasingly used in CSE, it differs fundamentally from CSE in predicting system behavior. Traditional CSE follows a deductive approach, deriving mathematical equations from conservation laws and physical properties to describe system causality. Moreover, CSE is used to predict the behavior of complex systems that are difficult or impossible to study experimentally [85]. In contrast, ML is an inductive discipline that learns relationships, potentially non-causal, directly from data or existing numerical models [107]. SciML is a subset of ML and is an interdisciplinary field aimed at accelerating scientific and engineering discoveries. While ML was once used only to help identify patterns in data from scientific instruments and sensors, SciML has evolved beyond data analysis to accelerate diverse scientific tasks, as discussed in the next section.
2.1 Scope
This paper focuses on SciML that enhances computational modeling and simulation of physical systems. Specifically, we address SciML approaches that support computational workflows where a model must be evaluated repeatedly within larger algorithmic procedures. In scientific computing, these are known as “outer-loop” processes because they form an additional computational layer that wraps around the core model solution. Such processes include: (1) explanatory modeling, which searches for insights into complex system behaviors; (2) uncertainty quantification, which infers model parameters from observations and propagates them to predictive distributions [28]; (3) sensitivity analysis, which identifies factors with greatest influence on predictions [88]; (4) experimental design, which predicts which data collections maximize understanding or decision value [87]; and (5) optimal design and control, which optimizes objective functions by adjusting decision variables under constraints [65, Ch. 1]. To support these outer-loop processes, predictive SciML encompasses four main classes of methods, each presenting different computational challenges: surrogate modeling, model discovery, hybrid CSE-SciML modeling, and outer-loop learning.
Surrogate modeling builds low-cost approximations of computationally expensive models for outer-loop analyses. Surrogate models can map model inputs to specific quantities of interest [22, 103, 10]. Surrogate models can also approximate complete solutions to governing equations via data-driven reduced-order modeling [81, 91, 12], operator learning with Gaussian processes [69] and neural networks [39, 54, 63, 8, 58, 76, 27], and physics-informed neural networks (PINNs) [114, 82, 30], which encode physical constraints directly into the neural network training. Finally, surrogate models can provide solutions to differential equations without simulation data [99, 24].
Model discovery learns equations approximating physical system principles [91, 12, 59, 18]. Like operator learning and reduced-order modeling, these approaches work with both experimental and simulation data. Model discovery is being increasingly used to develop constitutive closure models that can be used when a numerical model cannot practically resolve small scales such that the large-scale equations are unclosed.
Hybrid CSE-SciML models embed SciML components within CSE models. Examples include learning initial guesses for iterative PDE solvers [41], predicting preconditioners for linear systems [90], and developing constitutive models for applications ranging from radiation transport [42] to aerospace [98].
Outer-loop learning leverages ML to predict solutions under varying conditions. Applications include surrogate models for expensive forward simulations in PDE-constrained optimization [110, 64] and Bayesian inverse problems [14]. Reinforcement learning solves PDE-constrained optimization across different conditions [101]. ML methods can also learn solutions to these problems directly. For example, ML models learn Bayesian posterior distributions as marginals of joint distributions for fast online inference with arbitrary observational data [6, 7, 111].
2.2 Comparing CSE and SciML model development and deployment
Making scientific claims using SciML models involves two steps: model development and model deployment. While SciML models can complement CSE models, their development processes differ, though deployment processes remain similar, as shown in Figure 1. To illustrate these steps, we use an example quantifying uncertainty in a hybrid SciML-CSE ice-sheet model of the Humboldt Glacier in Greenland, with unknown glacier-landmass friction. Our example is loosely based on the study by He et al. [37], which exemplifies thorough reporting of a SciML model’s strengths and weaknesses.
CSE model development begins with physics-based governing equations. In our example, while nonlinear Stokes equations accurately model shear-thinning behavior for ice sheets, simpler models like shallow-shelf approximation (SSA) are often used for computational tractability [37]. The second step implements numerical methods to solve the governing equations while respecting properties like mass and energy conservation. This involves choosing a fixed ansatz (e.g., a finite element basis) and applying numerical procedures using linear algebra, Newton iteration, and time-stepping. For instance, He and coworkers [37] used continuous piecewise linear elements for thickness and velocity fields with Newton iteration and semi-implicit time-stepping.
In contrast to CSE, typical SciML development relies on training procedures that require observational data or the ability to query an existing numerical model. Consequently, as shown in Figure 1, the development begins with collecting input-output training data. For example, He et al. [37] built a DeepONet [63] to map ice-sheet thickness and friction to glacier velocity, using data from an SSA-based CSE model to accelerate glacier evolution and uncertainty quantification. The second step sets model hyperparameters (e.g., neural-network width and depth). The third optimizes model parameters (e.g., neural-network weights and biases) by minimizing the difference between model outputs and training data. He et al. [37] used Adam optimization to minimize the mean-squared error between predicted and training velocities. Often, hyperparameters must be tuned through formal optimization or informal assessment of model accuracy across different settings. Note that some SciML models [99, 24] combine CSE and SciML development steps shown in Figure 1, training without data by minimizing objectives that enforce known governing equations.
The development and deployment of SciML models for specific scientific applications follows distinct phases. When using these models to make scientific claims—–like predicting the Humboldt glacier’s impact on sea level in the future—–deployment involves three key steps beyond the initial model development. First, we collect calibration data representing the system’s current state (distinct from training data, which spans multiple states). Second, we calibrate the model through optimization to match this system-specific data. In our ice-sheet example, this means solving an inverse problem for friction parameters to match recent velocity measurements [46], notably optimizing model inputs rather than SciML parameters like neural-network weights. Finally, we run the model under various future conditions or plausible friction fields, such as those obtained from Bayesian calibration.
2.3 Leveraging an existing CSE framework for trustworthy SciML
As highlighted by Broderick et al. [11], trust in SciML and CSE models can be gained or lost at any step of development and deployment (Figure 1). Building trust requires demonstrating that the mathematical model adequately meets analysis goals despite being an approximation. The model must then be verified by comparing theoretical and empirical results to ensure it solves its mathematical formulation correctly. Additionally, rigorous code testing must be implemented to reduce software bugs, and data provenance must be documented to ensure relevance for future conclusions and decisions.
Over decades, the scientific computing community developed rigorous standards for building credibility in computational models. These standards crystallized into a systematic approach known as verification and validation (V&V), providing guidelines to enhance model trustworthiness by addressing potential failure points [1, 4, 72]. Verification ensures correct algorithm and code implementation (solving the equations correctly), while validation assesses model appropriateness for intended applications (solving the correct equations).
While V&V practices are well-established in scientific computing, the emergence of SciML brings new concerns that existing guidelines don’t fully address. SciML combines aspects of both scientific computing and ML, presenting unique verification and validation challenges beyond either field. We argue that adapting and extending proven V&V principles from scientific computing can provide a foundation for building trustworthy SciML models. To this end, we make recommendations in four areas: problem definition, verification, validation, and continuous credibility building (see Figure 2). Problem definition and verification (Sections 3 and 4) expand on the model development steps in Figure 1, while validation and reporting (Sections 5 and 6) detail the deployment steps. Validation is generally most pertinent only to papers that utilize SciML to support a scientific claim for a specific application. However, all other components illustrated in Figure 1 are applicable to all SciML papers, including those that present algorithmic advances, which are relevant to a broad spectrum of scientific applications.
Like Broderick et al. [11], we view trustworthiness as a spectrum rather than binary. Following Oberkampf’s predictive capability maturity model [75], the four components in Figure 2 build confidence through continuous testing and refinement as new data emerge. Each study should mature progressively in its development process through these four areas. The following sections identify opportunities for improving SciML practice, along with corresponding recommendations.
We do not claim that these four components or our recommendations are comprehensive or universally applicable. Instead, we aim to stimulate community efforts to develop SciML guidelines and standards. We hope to advance practice beyond the basic implementation of our recommendations toward more thorough, deliberate, and well-documented approaches to SciML development and deployment.
3 Problem definition
The first step in building and using a CSE or SciML model is defining the problem scope: the model’s intended purpose, application domain and operating environment, required quantities of interest (QoI) and their scales, and how prior knowledge informs model conceptualization.
3.1 Model purpose
A SciML model’s purpose, as discussed in Section 2.1, dictates all subsequent modeling choices. This purpose determines required outer-loop processes and essential properties.
To highlight the importance of specifying the target outer-loop process, consider a model used for explanatory modeling. An explanatory model must simulate all system processes, like ice-sheet thickness and velocity evolution. In contrast, a risk assessment model needs only decision-relevant quantities, like ice-sheet mass loss under varying emissions scenarios. Design and control models, meanwhile, have different requirements than those for risk assessment. The model purpose dictates the data types and formulations needed to train a SciML model. The impact of this purpose on data requirements and physics-informed objectives varies by application domain. Thus, the exact model formulation should be chosen in light of these problem-specific considerations.
3.2 Verification, calibration, validation and application domain
The trustworthy development and deployment of a model (see Figure 2) requires using the model in regards to verification, calibration, validation, and application domains. These domains are defined by the conditions under which the model operates during these respective phases and must be defined before model construction because they determine viable model classes. Key features include boundary conditions, forcing functions, geometry, and timescales. For ice sheets, examples include surface mass balance, land mass topography, and ocean temperatures.
Each domain will often require the prediction of different quantities of interest under different conditions. Moreover the complexity of the processes being modeled typically increase when transitioning from verification to calibration, to validation to prediction. Additionally the amount of data to complement or inform the model decreases as we move through these domains. For example, the verification domain for our ice-sheet examplar predicts the entire state of the ice-sheet for simple manufactured or analytical solutions. The calibration domain predicts Humboldt Glacier surface velocity under steady-state preindustrial conditions. The validation domain predicts grounding-line change rates from the first decade of this century. The application domain predicts glacier mass change in 2100. Figure 3 illustrates these distinct domains. When transitioning between domains, data shifts across domains must be considered. For example, a model trained only on calibration data from recent ice-sheet forcings may fail to predict ice-sheet properties under different future conditions.
3.3 Quantities of interest
Quantities of interest (QoI) are the model outputs required by users. Their form and scale depend on modeling purpose and application domain. We now discuss key considerations for QoI selection.
Risk assessment requires reproducing only decision-critical QoI. For ice-sheets, these include sea-level rise from mass loss and infrastructure damage costs. Design applications similarly need few QoI to evaluate objectives and constraints, like thermal and structural stresses in aerospace vehicles. Design models need accurate QoI predictions only along optimizer trajectories222For each design iteration the model may still need to be accurate across all uncertain model inputs, while risk assessment models must predict across many conditions. Explanatory modeling demands more extensive QoI sets, such as complete ice-sheet depth and velocity fields for studying calving. Therefore, simple surrogates often suffice for risk assessment and design, but explanatory modeling may require operators or reduced order models.
3.4 Model conceptualization
Model conceptualization, which follows problem definition, involves selecting model structure based on prior knowledge. While essential to CSE model development [45], this step requires clear identification of the application domain and relevant QoI.
Model structure selection draws on key prior knowledge: conservation laws, system invariances like rotational and translational symmetries. These guide method selection—for example, symplectic time integrators [86] preserve system dynamics properties. Moreover, this knowledge informs and justifies the selection of candidate model structures. A CSE modeler chooses between model types like lumped versus distributed PDE models, and linear versus nonlinear PDEs. The optimal choice depends on application domain, QoI, and available resources. For example, linear PDEs may introduce more error but their lower computational cost enables better error and uncertainty characterization for tasks like optimal design. Similar considerations guide SciML model selection. For example, Gaussian processes excel at predicting scalar QoI with few inputs and limited data, but become intractable for larger datasets without variational inference approximations [61]. In contrast, deep neural networks handle high-dimensional data but require large datasets. The intended use also shapes model structure and training, e.g., optimization applications require controlling derivative errors [10, 76] to ensure convergence [14, 64]. These approximation errors must be understood and quantified where possible.
CSE has a strong history of using prior knowledge to formulate governing equations for complex phenomena and deriving numerical methods that respect important physical properties. However, all models are approximate and the best model must balance usability, reliability, and feasibility [34]. While SciML methods can be usable and feasible, more attention is needed to establish their trustworthiness. In the following two sections, we discuss how CSE V&V can improve the trustworthiness of SciML research.
4 Verification
Verification increases the trustworthiness of numerical models by demonstrating that the numerical method can adequately solve the equations of the desired mathematical model and the code correctly implements the algorithm. Verification consists of code verification and solution verification, which enhance credibility and trust in the model’s predictions. Code and solution verification are well-established in CSE to reduce algorithmic errors. However, verification for SciML models has received less attention due to the field’s young age and unique challenges. Moreoever, because SciML models heavily rely on data, unlike CSE models, existing CSE verification notions need to be adapted for SciML.
4.1 Code verification
Code verification ensures that a computer code correctly implements the intended mathematical model. For CSE models, this involves confirming that numerical methods and algorithms are free from programming errors (“bugs”). PDE-based CSE models commonly use the method of manufactured solutions (MMS) to verify code on arbitrarily complex solutions. MMS substitutes a user-provided solution into the governing equations, then differentiates it to obtain the exact forcing function and boundary conditions. These solutions check if the code produces the known theoretical convergence rate as the numerical discretization is refined. If the observed order of convergence is less than theoretical, causes such as software bugs, insufficient mesh refinement, or singularities and discontinuities affecting convergence must be identified.
Code verification for SciML models is important but challenging due to the large role of data and nonconvex numerical optimization. Three main challenges limit code verification for many SciML models. First, while theoretical analysis of SciML models is increasing [93, 94, 78, 57, 55, 53, 54], many models like neural networks do not generally admit known convergence rates outside specific map classes [93, 94, 78, 38], despite their universal approximation properties [40, 20, 57]. Second, generalizable procedures to refine models, such as consistently refining neural-network width and depth as data increases, do not exist. Finally, regardless of data amount and model unknowns, modeling error often plateaus at a much higher level than machine precision due to nonconvex optimization issues like local minima and saddle points [21, 9].
Developing theoretical and algorithmic advances to address the three main challenges limiting code verification can substantially improve the trustworthiness of SciML models. Convergence-based code verification is currently possible only for certain SciML models with theory that bounds approximation errors in terms of model complexity and training data amount, such as operator methods [105], polynomial chaos expansions [16, 112], and Gaussian processes [13].
For SciML models without supporting theory, convergence tests should still be conducted and reported. Studies providing evidence of model convergence engender greater trustworthiness than those that do not, even when the empirically estimated convergence rate cannot be compared to theoretical rates. For example, observing Monte Carlo-type sampling rates in a regime of interest for a fixed overparametrized model can provide intuition into whether the model should be enhanced.
To account for the heavy reliance of SciML models on training data optimization, code verification should be adapted in two ways. First, report errors in the ML model for a given complexity and data amount for different realizations of the training data to quantify the impact of sampling error, which is not present in CSE models. Second, because most SciML algorithms introduce optimization error, conduct verification studies that artificially generate data from a random realization of an ML model, then compare the recovered parameter values with the true parameter values or compare the predictions of the learned and true approximations, or at the very least compare the predictions of the two models. Additionally, quantify the sensitivity of the SciML model error to randomness in the optimization by varying the random seed and initial guess passed to the optimizer (see Section 6.3). All verification tests must employ test data or verification data, independent of the training data, to measure the accuracy of the SciML model.
4.2 Solution verification
Code verification establishes a code’s ability to reproduce known idealized solutions, while solution verification, performed after code verification, assesses the code’s accuracy on more complex yet tractable problems defined by more realistic boundary conditions, forcing, and data. For example, code verification of ice sheets may use manufactured solutions, whereas solution verification may use more realistic MISMIP benchmarks [17]. In solution verification, the numerical solution cannot be compared to a known exact solution, and the convergence rate to a known solution cannot be established. Instead, solution verification must use other procedures to estimate the error introduced by the numerical discretization.
Solution verification establishes whether the exact conditions of a model result in the expected theoretical convergence rate or if unexpected features like shocks or singularities prevent it. The most common approach for CSE models compares the difference between consecutive solutions as the numerical discretization is refined and uses Richardson extrapolation to estimate errors. A posteriori error estimation techniques that require solving an adjoint equation can also be used.
While thorough solution verification of CSE models is challenging, these difficulties are further amplified for SciML models. Currently, solution verification of SciML models simply consists of evaluating a trained model’s performance using test data separate from the training data. However, this is insufficient as solution verification requires quantifying the impact of increasing data and model complexity on model error. Yet, unfortunately, performing a posteriori error estimation for many SciML models using techniques like Richardson extrapolation is difficult due to the confounding of model expressivity, statistical sampling errors, and variability introduced by converging to local solutions or saddle points of nonconvex optimization, making it challenging to monotonically decrease the error of SciML models such as neural networks.
Until convergence theory for SciML models improves and automated procedures are developed to change SciML model hyperparameters as data increases, solution verification of SciML models should repeat the sensitivity tests proposed for code verification (Section 4.1) with two key differences: First, verification experiments used to generate verification data must be specifically designed for solution verification, as not all verification data equally informs solution verification efforts, similar to observations made when creating validation datasets for CSE models [72]. See Section 6.1 for more information on important properties of verification benchmarks. Second, while ideally the convergence of SciML errors on realistic benchmarks should be investigated, it may be computationally impractical. Thus, solution verification should prioritize quantifying errors using the model complexity and data amount that will be used when deploying the SciML model to its application domain.
5 Validation
Verification establishes if a model can accurately produce the behavior of a system described by governing equations. In contrast, validation assesses whether a CSE model’s governing equations—or data for SciML models—and the model’s implementation can reproduce the physical system’s important properties, as determined by the model’s purpose.
Validation requires three main steps: (1) solve an inverse problem to calibrate the model to observational data; (2) compare the model’s output with observational data collected explicitly for validation; and (3) quantify the uncertainty in model predictions when interpolating or extrapolating from the validation domain to the application domain. We will expand on these steps below. But first note that the issues affecting the verification of SciML models also affect calibration and validation. Consequently, we will not revisit them here but rather will highlight the unique challenges in validating SciML models.
5.1 Calibration
Once a CSE model has been verified, it must be calibrated to match experimental data that contains observational noise. This calibration requires solving an inverse problem [100], which can be either deterministic or statistical (e.g., Bayesian). The deterministic approach formulates the inverse problem as a (nonlinear) optimization problem that minimizes the mismatch between model and experimental data. This formulation requires regularization to ensure well-posedness, typically chosen using the L-curve [35] or the Morozov discrepancy principle [3]. The Bayesian approach replaces the misfit with a likelihood function based on the noise model, while using a prior distribution for regularization. This prior distribution ensures well-posedness while encoding typical parameter ranges and correlation lengths. We recommend Bayesian methods for calibration because they provide insight into the uncertainty of the reconstructed model parameters.
The calibration of SciML operator, reduced-order, and hybrid CSE-SciML models is distinct from SciML training and follows similar principles to CSE model calibration. These models are first trained using simulation data for solution verification. Next, observational data (called calibration data) determines the optimal model input values that match experimental outputs. For instance, calibrating a SciML ice-sheet model such as that of Ref. [37] requires finding optimal friction field parameters of a trained SciML model, which best predict observed glacier surface velocities, given the noise in the observational data.
Calibration typically improves a model’s predictive accuracy on its application domain, but the informative value of calibration data varies significantly. Therefore, researchers should select calibration data strategically to maximize information content within their experimental budget. See Section 6.1 for further discussion on collecting informative data.
5.2 Model validation
Model validation is the “substantiation that a model within its domain of applicability possesses a satisfactory range of accuracy consistent with the intended application of the model” [83]. Validation involves comparing computational results with observational data, then determining if the agreement meets the model’s intended purpose [56]. For CSE models with unacceptable validation agreement, modelers must either collect additional calibration data or refine the model structure until reaching acceptable accuracy. SciML models follow a similar iterative process but offer an additional option: to collect more training data.
Model validation must occur after calibration and requires independent data not used for calibration or training. For our conceptual ice-sheet model, calibration matches surface velocities assumed to represent pre-industrial conditions, while validation assesses the calibrated model’s ability to predict grounding line change rates at the start of this decade. Performance metrics must target the specific modeling purpose. For optimization tasks, metrics should measure the distance from true optima obtained via the SciML model or bound the associated error [14]. For uncertainty estimation, metrics should quantify errors in uncertainty statistics through moment discrepancies or density-based measures like (shifted) reverse and forward Kullback–Leibler divergences. For explanatory SciML modeling, validation metrics must also assess physical fidelity: adherence to physical laws, conservation properties (such as mass and energy), and other constraints. As with verification, the validation concept should encompass data validation, particularly whether training data adequately represents the application space.
Validation determines whether a model is acceptable for its specific purpose rather than universally correct. The definition of acceptable is subjective, depending on validation metrics and accuracy requirements established by model stakeholders in alignment with the problem definition and model purpose (see Section 3). Moreoever, validation itself does not constitute final model acceptance, which must be based on model accuracy in the application domain, as discussed in Section 5.3.
Two additional considerations complete our discussion of model validation. First, this validation differs from the concept of cross validation, which estimates ML model accuracy on data representative of the training domain during development. The validation described here assesses accuracy in a separate validation domain. Second, validation data varies in informative value. Validation experiments should “capture the essential physics of interest, including all relevant physical modeling data and initial and boundary conditions required by the code” [73]. Most critically, validation data must remain independent from training and calibration data.
5.3 Prediction
Although extensive data may be available for model calibration, validation data is typically scarcer and may not represent the model’s intended application domain. According to Schwer [95], “The original reason for developing a model was to make predictions for applications of the model where no experimental data could, or would, be obtained.” Therefore, minimizing validation metrics at nominal conditions cannot sufficiently validate a model. Modelers must also quantify accuracy and uncertainty when predictions are extrapolated to the application domain.
SciML models, like CSE models, are subject to numerous sources of uncertainty. The impact of these uncertainties on model predictions must be quantified. Several sources of uncertainty affect CSE models. These include: numerical errors, from approximating the solution to governing equations; input uncertainty arises, which is caused by inexact knowledge of model inputs; parameter uncertainty, which stems from inexact knowledge of model coefficients; and model structure error representing the difference between the model and reality. SciML models contain all these uncertainties. They also incorporate additional uncertainties from sampling and optimization errors, as discussed previously.
Sampling error arises from training a model with a finite amount of possibly noisy data. For a fixed ML model structure and zero optimization error, this error decreases as the amount of data increases. Optimization error represents the difference between the optimized solution, which is often a local optimum, and the global solution for fixed training data. Optimization error can enter CSE models during calibration. Optimization error affects SciML models more significantly because it occurs both during calibration and training. Linear approximations, for example, based on polynomials, achieve zero optimization error during training to machine precision. However, nonlinear approximations such as neural networks often produce non-trivial optimization errors. Stochastic gradient descent demonstrates this by producing different parameter estimates due to stochastic optimization randomness and initial guesses.
The identified sources of modeling uncertainty require parameterization for sampling. Expert knowledge typically guides the construction of prior distributions that represent parametric uncertainty. This parameterization should occur during problem definition. Bayesian calibration updates these priors into posterior distributions using calibration data. The model must then propagate all uncertainties onto predictions in the application domain. Monte Carlo quadrature accomplishes this propagation by drawing random samples from the uncertainty distributions. The method collects model predictions at these samples and computes empirical estimates of important statistics defined by validation criteria, such as mean and variance.
We emphasize the impact of all sources of error and uncertainty must be quantified. Simply estimating the impact of error caused by using finite sample sets, for example estimated by generative models such as variational autoencoders of Gaussian processes is insufficient. Moreover, complete elimination of uncertainty is impossible. Consequently, model acceptance, like validation, must rely on subjective accuracy criteria established through stakeholder communication. For instance, acceptance criteria for predicted sea-level change from melting ice-sheets at year 2100 may specify that prediction precision reaches 1% of the mean value. Yet, engineering applications, such as those focused on aerospace design, may have much higher accuracy requirements.
The aforementioned Monte-Carlo based UQ procedure effectively quantifies the impact of parameterized uncertainties on model predictions. However, model structure error remains difficult to parameterize in both SciML and CSE modeling. Validation can partially assess model structure error. However, experiments rarely cover all conditions of use. Specifically, validation tests only the model’s interpolation ability within the convex hull of available data and assumptions. This limitation creates challenges when applying the model outside its validation domain. Some progress exists in quantifying extrapolation error for “models based upon highly-reliable theory that is augmented with less-reliable embedded models” [77]. However, such hybrid CSE-SciML models rely on well-established physics-based governing equations to support extrapolation confidence. Pure SciML models still require substantial research to develop reliable methods for estimating model structure uncertainty.
6 Continuous Credibility Building
The three model development components depicted in the top of Figure 2 appear sequential. However, practical SciML development is iterative, often requiring steps to be repeated. Poor validation metrics may necessitate model updates or additional data collection. Sometimes, simplifying objectives or the QoI scale may be warranted to achieve more accurate results for a simpler but still valuable problem.
The iterative nature of modeling requires ongoing documentation of the critical aspects that affect V&V including data sources and processing, loss functions and optimizers and model tuning. Additionally, it is important to perform software quality assurance (SQA), model comparisons, and explain and interpret predictions and model behaviors. Enacting these steps is referred to as continuous credibility building in Figure 2.
6.1 Data Sources
In addition to listing prior knowledge that can be used to build an ML model, the types of data used to train a model affects model V&V. The data sources, which will be used to train the SciML model, must be carefully chosen and the reasons for those choices must be documented. Important data properties to consider are data fidelity and bias, data noise and the cost of collecting training data, which limits the amount that can be collected. Some studies make recommendations on the amount of data that should be used to train an ML model based on the number of model input features [117]. However, we believe the amount of data should be defined by the accuracy requirements of the intended use of the SciML model. Future work on generalization accuracy and statistical learning theory for SciML models will help better elucidate these mathematical trade-offs.
When training a SciML model, the fidelity of data impacts the performance of a model. For example, when training using simulation data, the numerical discretization of the model generating the data is important. Ideally, training data should be generated using the most accurate (highest-fidelity) discretization. However, often a cheaper but less accurate discretization (lower-fidelity) is used because it allows for more training data to be collected. While more low-fidelity training data allows a better approximation of the low-fidelity simulation model, it introduces a bias in the SciML model relative to the high-fidelity model. The impact of this error on quantities predicted by a model, for instance the optimal objective of a design problem, must be quantified. Alternatively, multi-fidelity SciML models, which combine limited high-fidelity data with larger amounts of low-fidelity data [29, 67, 19], can sometimes be used to mitigate this issue.
The creation of informative benchmarks can substantially improve the quality of V&V analyses and enable easier comparison of methods and SciML structures. However, unlike traditional ML, SciML, and CSE, lack a set of commonly-used and agreed-upon benchmark problems, for both solution verification and validation. Consequently, further work is needed to complement early efforts to create SciML benchmarks [102]. Motivated by the work of Oberkampf [74], we believe the purpose and details of the benchmark should be precisely stated, including how comparisons with the benchmark must be made and the the exact verification or validation acceptance criteria. For example, a natural acceptance criteria for a SciML surrogate used for UQ is that the surrogate can accurately compute the mean and variance of the true model it replaces to a specified level of accuracy. In contrast, a reasonable acceptance criterion of a SciML model used for optimization would be based on the comparison of the optimal solution obtained by the SciML model with known or trusted reference optima. Lastly, benchmarks must be made widely accessible, though this can be difficult when some benchmarks are associated with terabytes of data. Thus, until community benchmarks and mechanisms for their dissemination are developed, it is crucial that researchers thoroughly document these choices and explain the significance of the problems they create.
6.2 Data processing
Before an ML model can be constructed, the data that will be used to train the ML model must be processed for missing and abnormal data.333It goes without saying that unprocessed raw data should be preserved and properly archived before any transformations are applied. When building ML models from simulation data, missing data can occur when one or more simulations fail to complete, caused for instance by the model becoming unstable for a certain set of inputs, or the computer system crashing. Similarly, abnormal data can be generated when the instability of the model does not cause the simulation to fail but leads to errors, or when a soft fault occurs on a high-performance computing system. The methods employed for dealing with abnormal data—such as removal, interpolation or imputation—must be defined as they can substantially impact the accuracy and credibility of a SciML model.
Variable transformations are also often used to process data when training a SciML model. For instance, it is common to normalize the data to have zero mean and unit standard deviation, or to take the logarithm of the data with the goal of improving the predictive performance of a SciML model. Modelers should report the use of any transformations because the use of different transformations can lead to substantial changes to the predictive performance of the ML model. Moreover, data processing scripts should be preserved in a trusted repository and published alongside other study artifacts.
6.3 Loss functions and optimizers
Training an ML model requires minimizing a loss function that quantifies how well the model approximates the training data. This loss function is often the mean squared error or mean absolute error. Nevertheless, a plethora of other loss functions can be used. Indeed, the loss function should be tailored to the intended use of the ML model. For example, the mean squared error captures the average size of the squared residuals between the ML model and the data, but treats positive and negative residuals equally. Thus, if learning a function intended to capture large events it may be more appropriate to train the SciML model using asymmetric loss functions such as the quantile loss function [47].
Additionally, the optimization method used to minimize the loss function can also substantially impact model trustworthiness. For example, it is common to use stochastic optimization methods such as Adam [52] to train a SciML model. Thus, the optimal solution depends on the random seed, the initial condition and the computing platform used to perform the optimization. Consequently, the impact of this randomness must be quantified, for example, by reporting the variability in the error of the SciML model learned using different perturbations of these factors.
6.4 Model tuning
Training a SciML model is typically an iterative process, which involves optimizing the parameters of a model, e.g., weights and biases of a neural network, for different values of the model hyper-parameters, e.g., the depth and width of a neural network, also called hyperparameter tuning [26]. Model structures with large numbers of parameters (common in SciML) can lead to overfitting, where the model performs well on the training data but poorly on new, unseen data. Additionally, an overly complex model can become less interpretable, making it harder for researchers to understand and explain the model’s predictions. Furthermore, it typically becomes harder to identify the impact of each parameter, a problem known as reduced parameter identifiability [31], as the number of model parameters increases. The final choice of hyperparameters can significantly impact a model’s predictive power, interpretability, and the uncertainty of its predictions. Yet, despite the importance of hyperparameter tuning, a review of over 140 highly cited papers using ML for environmental research [117] found that only 19.6% of those papers conducted formal hyperparameter optimization methods (such as grid search) and another 25.7% used trial and error. To ensure that the scientific community can fully evaluate and understand published works in SciML, detailed documentation of these processes is essential.
6.5 Software quality assurance and reproducibility
Software quality assurance (SQA) is required by both CSE and SciML models to minimize the impact of programming errors in model predictions, facilitate collaboration, and build credibility into the modeling activity. Some CSE guidelines call for SQA as part of code verification [72], given its focus on correctness of the software by finding and fixing programming errors or ‘bugs’ that are prevalent in all software [36, 115]. While SQA is an essential component of code verification, we discuss it separately here to emphasize its importance to the entirety of CSE and SciML workflows.
SQA requires both static analysis [62, 25] and dynamic testing [48]. Static analysis checks for errors without executing code and can be applied to an entire code base before dynamic tests are executed. This is useful because runtime tools, such as Valgrind [71], only test the code that is executed and thus can miss corner cases that are not tested [32]. Dynamic testing comprises unit tests, integration tests and system tests. All of these software engineering practices have low adoption in ML software [97]. Software engineering is particularly challenging in SciML models because it has “…all of the problems of non-ML software systems plus an additional set of ML-specific issues” [108]. While deterministic CSE can often use theoretically derived test “oracles” that capture the expected behavior of a model [49], ML’s stochastic nature requires aggregating multiple model runs to assess variability. System-wide testing is often needed to understand emergent properties, making unit testing less useful. Additionally, code coverage metrics are less meaningful for SciML since decision logic is relearned with each training [116]. Despite the importance of SQA, it is often neglected in both CSE and SciML. An analysis of 62 scientific software studies found that most used only one type of testing, and none used all types (unit, system, integration, acceptance, and regression testing) [50]. Improving SQA adoption is thus crucial for building trustworthy SciML models. SciML studies should make every effort to establish confidence in the software implementation, via modular tests and benchmarks, and adoption of software engineering best practices.
Reproducibility is essential for credibility building of computational models [70], and can be understood as another form of quality assurance. For reproducibility, authors should provide comprehensive documentation, including data specifications, computing requirements, software versions, and public code repositories. The optimized parameter and hyper-parameters of SciML models, e.g., the width and depth of a neural network and its weights and biases, should also be released. Repositories should be well-organized with tutorials and reproduction scripts [51], and standard public licenses should be adopted that stipulate reuse rights of the artifacts.
Warnings about lack of reproducibility in ML research have been raised before [43, 33]. A common failing, with an obvious solution, is not sharing the source code associated with the published research. Yet simply depositing the software in a public repository is far from sufficient to ensure reproducibility in ML research. Rigorous documentation of the training pipeline is necessary, including details of the optimization and regularization applied, activation functions, hyperparameters, learning-rate schedules, and more. Data provenance, as discussed earlier, should be managed strictly and documented; ideally the full data sets should be deposited in trusted repositories that provide a persistent identifier. None of these efforts is minor, and sometimes they may seem unrewarding, yet they are necessary, and increasingly paid attention to by funders, journals, and conferences. Higher reproducibility standards will advance SciML’s trustworthiness, and benefit authors by reduced barriers to the adoption of their proposed methods.
6.6 Cost-accuracy comparison with alternatives
SciML models must be compared with state-of-the-art alternatives throughout model development, including verification, validation, and prediction stages. However, a review of environmental SciML studies found that 37.8% made no ML comparisons, 44.6% compared with one SciML method, and only 14.9% compared with multiple methods [117].
Model comparison credibility depends on the types of SciML models compared. When a parsimonious model outperforms more complex ones [113], comparing it only within its model family is insufficient. Comparing a neural network with Gaussian processes and linear regression better contextualizes performance than comparing multiple neural networks. While comparing multiple ML methods is valuable, when CSE alternatives exist, SciML methods must also be evaluated against state-of-the-art CSE approaches using purpose-specific metrics, particularly those used in validation.
Although accuracy is crucial in method comparison, computational costs must also be documented, yet total modeling costs are often underreported. Researchers should detail all costs, including data generation (e.g., running a CSE model), parameter optimization, and prediction. They should also compare the total of these costs against state-of-the-art alternatives. For instance, in PDE-constrained optimization, direct PDE optimization costs should be compared to combined surrogate training and optimization costs.
Dismissing training costs as “offline” is misleading, however, these costs can be amortized when SciML models are used across multiple studies, as in learning to optimize [15]. Nevertheless, such amortization claims need qualification since SciML models’ cost-effectiveness—the trade-off between accuracy, development time, and cost—is harder to predict than for CSE models [96].
Finally, every effort must be made to avoid “the misuse of language, intentionally or unintentionally, as this affects the interpretation of study findings and has been described as “spin” [2] or reporting bias. While this may appear obvious, a recent review of 152 studies of clinical prediction using ML models found that 95% of abstracts that recommended the use of an ML in daily practice lacked any external validation of the developed models. Moreover, only 55.6% of studies recommended the clinical use of an ML model in their main text without any external validation. McGreivy and Hakim [66] found that 79% of 76 ML papers solving fluid PDEs used weak baselines when claiming to outperform standard methods. The same study also found that outcome reporting bias and publication bias are widespread. Of 232 abstracts surveyed, 94.8% reported only positive results, 5.2% reported mixed results, and none reported only negative results, suggesting reporting bias.
6.7 Explain and Interpret
SciML and CSE researchers should explicitly identify how their methods advance specific scientific objectives. Sufficient information should be provided “to infer the purpose of the technology and set their expectations sensibly based on such understanding” [104]. Model trustworthiness depends both on the process and its results. Consequently, researchers should explain how they came to their conclusions. According to Doshi-Velez and Kim [23], this can involve experts interpreting highly intelligible models, where interpretability is built into their model structure. For example, a physicist might understand the reasons for the bifurcation of a dynamical system or use a second model to understand the learned model’s behavior. Doshi-Velez and Kim [23] also advocate for evaluating algorithm robustness to parameter and input variations and for validating that predicted changes in model outputs correspond to real system behavior.
The limited interpretability and transparency of ML models are significant challenges, especially in critical scientific applications such as medicine and aerospace engineering. The inner workings of SciML models are hard for humans to understand, possibly deterring us from trusting their predictions. This is especially problematic in settings where model-assisted decisions can significantly impact people’s lives. To address these problems, researchers must continue to develop new techniques for making SciML models more interpretable and transparent.
While many papers compare the interpretability of different methods, no universal definition of interpretability exists because interpretability is subjective and varies across scientific disciplines [60]. Linear models are often considered interpretable because coefficients show each feature’s importance in predictions. However, this interpretability breaks down with thousands of parameters [68]. Claims of superior interpretability must be thoroughly justified. This creates an opportunity for research communities to define domain-specific properties that make models interpretable, complemented by explainability methods that reveal how models make their predictions.
7 Concluding remarks
Scientific machine learning (SciML) is an emerging field that integrates machine learning into scientific workflows, creating a powerful synergy between traditional and new methodologies. More specifically, we define SciML as the application of machine learning methods to enhance, accelerate, or improve the computational modeling and simulation of physical systems. This definition is applicable across a wide array of scientific disciplines, including physics, earth sciences, and engineering. However, we acknowledge that this definition may not be universally embraced in certain fields, such as psychology and social sciences.
While computational science and engineering (CSE) has developed solid theoretical foundations over decades, SciML’s rapid empirical advances have often outpaced formal theoretical analysis. Successfully advancing SciML requires thoughtfully integrating the complementary strengths of both fields: CSE’s mathematical rigor and ML’s data-driven innovations. Moreover, the CSE community has established rigorous guidelines for model development, verification, and validation (V&V) to support scientific claims. In contrast, such standards are still emerging in SciML. Consequently, this paper addresses the critical need for developing V&V practices for SciML that builds upon CSE frameworks while addressing SciML’s unique challenges.
In this paper, we identified key challenges to developing trusted SciML practices, including the field’s heavy reliance on data, the optimization of non-convex functions, and gaps in theoretical understanding. To address these challenges, we presented a four-component framework for developing and deploying trustworthy SciML models. As part of this framework, we established 16 recommendations to provide concrete guidance on critical aspects like reporting computational requirements and preparing detailed documentation. We also used real-world examples to illustrate concepts and offer actionable guidance for both development and deployment phases.
The recommendations in this paper are particularly crucial for high-consequence applications where model credibility is paramount. These include safety-critical systems in aerospace and nuclear engineering, digital twins for infrastructure monitoring, and medical applications that use model predictions guide treatment decisions. In such domains, the cost of model failure—whether from inadequate verification, poor validation, or incomplete uncertainty quantification—can be severe. Our framework provides a foundation for building the necessary trust in SciML models for these critical applications.
Beyond these domain-specific challenges, the rapid evolution of the field and the drive to publish quickly have led to poor reproducibility, often compromising rigor and transparency. Here, we aim to catalyze broader community dialogue around establishing consensus-based practices for predictive SciML. As such, this paper makes a meaningful contribution by being one of the first comprehensive attempts to address systematic V&V practices specifically for SciML. It fills an important gap in the literature, as while both CSE and ML have their own established practices, the intersection of these fields requires new approaches to ensure trustworthiness.
We encourage journal editors to adopt these, or similar, recommendations in their submission guidelines, funding agencies to require V&V plans in proposals, conference organizers to promote sessions on SciML credibility, and researchers to implement these practices in their work. Only through collective effort can we establish rigorous standards needed to fully realize SciML’s potential for accelerating scientific discovery and enabling trustworthy predictions in high-consequence applications.
Acknowledgments
Funding: John Jakeman’s work was funded by Sandia National Laboratories’ Laboratory Directed Research and Development (LDRD) program. Thomas O’Leary-Roseberry’s work was funded by the National Science Foundation under DMS Award 2324643 and OAC Award 2313033. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan.
AI Disclosure: During the preparation of this work the authors used Sandia National Laboratories’ SandiaAI Chat and Claude.ai in order to correct spelling and grammar and reduce the length of the manuscript. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.